# Gaussian Eigen Models for Human Heads

### v3

@@ 1 Introduction
Half a century ago, Frederick Parke described a representation and animation technique to generate ,,animated sequences of a human face changing expressions” [54]. Using polygonal meshes, single facial expression states were described that could be combined with linear interpolation to generate new expression states (the ,,simplest way consistent with natural motion” [54]). Based on this principle, Blanz, and Vetter [3] introduced the so-called 3D morphable model (3DMM) - a statistical model of the 3D shape and appearance of human faces. Principle Component Analysis (PCA) is performed on a set of around 200 subjects that have been laser-scanned and registered to a consistent template to find the displacement vectors (principal components) of how faces change the most, in terms of geometry and albedo. With this PCA basis, new faces can be generated by specifying the coefficients for the principle components taking a dot product of the coefficients with the basis to obtain offsets, and adding them to the mean. State-of-the-art reports on face reconstruction and tracking [93], as well as on morphable models [8], state that this representation is widely used for facial performance capturing (regression-based and optimization-based) and builds the backbone of recent controllable photo-realistic 3D avatars that are equipped with neural rendering [68, 69, 87, 19, 12].
@@
Inspired by the simplicity of such mesh-based linear morphable models and addressing the lack of appearance realism of current 3DMMs, we propose a personalized linear appearance model based on 3D Gaussians as geometry primitive following 3D Gaussian Splatting (3DGS) [29]. In contrast to the work on Dynamic 3D Gaussian Avatars [41, 58, 91, 52, 85, 82, 64], our goal is a compact and light representation that does not need vast amounts of compute resources to generate novel expressions of the human. Unfortunately, most of the methods show that to produce high-quality results, one needs to employ heavy CNN-based architectures which are not well suited for commodity devices and tend to slow down the rendering pipeline. Moreover, those models comprise dozens of millions of parameters creating heavy checkpoints that can easily exceed 500 MB. This ultimately creates a major issue for distributing and managing personalized models. We tackle this problem by distilling a CNN-based architecture, leading to a personalized Gaussian Eigen Models for Human Heads, GEM in short. Our approach builds on Gaussian maps predicted from a modified UNet architecture [75] which is used for the UV space normalization required to properly build linear eigenbases. Based on the per-subject trained CNN model, we bootstrap the personalized GEM by computing an ensemble of linear bases on the predicted Gaussian maps of the training frames. The bases are refined on the training corpus using photometric losses while preserving their orthogonality.
@@
These lightweight appearance bases are controlled with a relatively low number of parameters ranging from twenty up to fifty coefficients which can be specified w.r.t. the available compute resources and can for example be regressed by a ResNet-based model [10]. We demonstrate this for the scenarios of self-reenactment as well as cross-person animation. We showcase a real-time application of our method in the supplemental video.
@@
    Gaussian Eigen Models for Human Heads (GEM) is a distillation technique of 3D Gaussian head avatar models built upon an ensemble of eigenbases.
@@ 2 Related Work
Figure 2: Given a multi-view video of a subject and mesh tracking, we create a dataset of 3D Gaussian point clouds for each frame in the sequence. Using this data, we distill a high-quality Gaussian Eigen Models for Human Heads (GEM). GEM is an ensemble of linear bases for each Gaussian primitive modality: position, opacity, scale, and rotation. Based on these bases, facial appearances are generated by a linear combination.
@@ 2.1 NeRF-based avatars
Numerous methods [91, 41, 64, 82, 86, 58, 11, 79, 25, 52, 34, 16, 92] capitalize on the speed and quality of 3DGS. Qian et al. [58] attach 3D Gaussians to the FLAME [40] mesh surface and apply a deformation gradient similar to Zielonka et al. [89] to orient the Gaussians according to the local Frenet frames of the surface. This method, however, does not utilize any information about expressions and, thus, struggles with pose-dependent changes (e.g., wrinkles, self-shadows) and, despite high-quality results, retrieves only a global static appearance model. 3D Gaussian blendshapes [46] controls an avatar by linearly interpolating between optimized blendshapes using 3DMM expression coefficents. However, this method depends on underlying 3DMM whereas GEM is a mesh-free representation. Li et al. [41] use a StyleUNet-like CNN architecture [75] to regress front and back Gaussian maps. Employing the powerful CNN network on position maps, they achieve impressive results for human bodies with effects like pose-dependent wrinkle formation.
@@ 3.1 Gaussian Eigen Model (GEM)
We are following the idea of organizing the 3D Gaussians in 2D maps [41, 52, 82, 64], where each pixel represents a 3D Gaussian with its parameters. We propose an adapted CNN-architecture of Animatable Gaussians (AG) [41], by merging the separate Style-U-Nets, reducing the convolutional layers, and operating in the UV space of the FLAME head model. In addition, we are employing deformation gradients following Sumner et al. [65] to handle the transformation from canonical to deformed space and treat the color as a global parameter. We refer to the suppl. mat. for a detailed explanation of the architectural changes. In comparison to the original AG model, our proposed CNN model produces slightly better results while being more efficient in terms of computing and memory. Using this model, we generate the per-frame Gaussian primitives $\boldsymbol{G}_{i}$ in the canonical space for all training time-frames. Note that for this reconstruction, we follow Animatble Gaussians and, thus, FLAME-based tracking is required. However, during inference, our model is independent of FLAME.
@@ 4.2 Cross-person Reenactment Evaluation
Figure 8: Compression error depending on the number of used principle components in GEM. The heatmaps show the photometric $\ell_{1}$-error for 10, 30, and 50 components using $128^{2}$ Gaussian maps. See suppl. doc. for additional evaluations.
Drawing inspiration from EMOCA [7], we further assess cross-re-enactment quantitatively by leveraging emotion recognition feature vectors from both the source image and the resulting cross-re-enactment, utilizing EmoNet [74]. For each pair of input and output images, we predict EmoNet features and measure cosine distance and $\mathcal{L}_{1}$ error between them. We report the numbers in the Table 3. Additionally, we also report FID scores [21] and rendering speed. Our method achieves on-par quality with a CNN-based solution while maintaining the highest frame rates and significantly outperforming GA in terms of quality.
@@ 4.3 GEM Ablation Studies
We are interested in the compression error introduced by the projection on different amounts of principal components used in GEM, also concerning the memory consumption. Additionally, we show that GEM is not topology-dependent. Figure 9 demonstrates that our method is able to handle different topologies (subject wearing glasses), despite utilizing a fixed UV space.
@@
For instance, our smallest model weighs as little as 7MB using only 10 components of the eigenbasis. This is almost 12 times less than our smallest CNN-based model and almost 70 times less than Animatable Gaussians [41]. In contrast to neural networks, we can easily trade quality over size which is very useful in the context of different commodity devices with reduced compute capabilities. Table 4 presents how compression affects the quality of reconstruction, where we evaluate a sequence with $\sim 1$k frames for a single actor under a novel view. As expected, using only 10 components impacts the quality the most, however, the results are still of high quality, see Figure 8. Gaussian Avatars [58] offers a small size of the stored Gaussians cloud, ranging from 5MB, and 14MB without the FLAME model for $128^{2}$, $256^{2}$ Gaussians, respectively. However, the quality of reconstruction lacks wrinkle details and sharpness as can be seen in Figure 13. In comparison to Gaussian Avatars [58], our model does not require FLAME during inference which is an additional 90MB.
@@ B.2 CNN Training Details
Figure 13: The quality comparison to Gaussian Avatars [58] shows better performance even though our model is similar in size to its Gaussian cloud and we do not need an additional FLAME model which weighs 90MB.

### v4

@@ 1 Introduction
Half a century ago, Frederick Parke described a representation and animation technique to generate ,,animated sequences of a human face changing expressions” [54]. Using polygonal meshes, single facial expression states were described that could be combined with linear interpolation to generate new expression states (the ,,simplest way consistent with natural motion” [54]). Based on this principle, Blanz, and Vetter [3] introduced the so-called 3D morphable model (3DMM) - a statistical model of the 3D shape and appearance of human faces. Principle Component Analysis (PCA) is performed on a set of around 200 subjects that have been laser-scanned and registered to a consistent template to find the displacement vectors (principal components) of how faces change the most, in terms of geometry and albedo. With this PCA basis, new faces can be generated by specifying the coefficients for the principle components taking a dot product of the coefficients with the basis to obtain offsets, and adding them to the mean. State-of-the-art reports on face reconstruction and tracking [93] as well as on morphable models [8] state that this representation is widely used for facial performance capturing (regression-based and optimization-based) and builds the backbone of recent controllable photo-realistic 3D avatars that are equipped with neural rendering [68, 69, 87, 19, 12].
@@
Inspired by the simplicity of such mesh-based linear morphable models and addressing the lack of appearance realism of current 3DMMs, we propose a personalized linear appearance model based on 3D Gaussians as geometry primitives following 3D Gaussian Splatting (3DGS) [29]. In contrast to the work on Dynamic 3D Gaussian Avatars [41, 58, 91, 52, 85, 82, 64], our goal is a compact and light representation that does not need vast amounts of compute resources to generate novel expressions of the human. Unfortunately, most of the methods show that to produce high-quality results, one needs to employ heavy CNN-based architectures which are not well suited for commodity devices and tend to slow down the rendering pipeline. Moreover, those models comprise dozens of millions of parameters creating heavy checkpoints that can easily exceed 500 MB. This ultimately creates a major issue for distributing and managing personalized models. We tackle this problem by distilling a CNN-based architecture, leading to a personalized Gaussian Eigen Models for Human Heads, GEM in short. Our approach builds on Gaussian maps predicted from a modified UNet architecture [75] which is used for the UV space normalization required to build linear eigenbases. Based on the per-subject trained CNN model, we bootstrap the GEM by computing an ensemble of linear bases on the predicted Gaussian maps of the training frames. The bases are refined on the training corpus using photometric losses while preserving their orthogonality.
@@
These lightweight appearance bases are controlled with a relatively low number of parameters ranging from twenty up to fifty coefficients which can be specified w.r.t. the available compute resources and can for example be regressed by a ResNet-based model [10]. We demonstrate this for self-reenactment as well as cross-person animation, including a real-time demo in the suppl. video.
@@
    Gaussian Eigen Models for Human Heads (GEM), a distillation technique of 3D Gaussian head avatar models built upon an ensemble of eigenbases.
@@ 2 Related Work
Figure 2: Given a multi-view video of a subject and mesh tracking, we create a dataset of 3D Gaussian point clouds for each frame in the sequence. Using this data, we distill a high-quality Gaussian Eigen Model (GEM). GEM is an ensemble of linear bases for each Gaussian primitive modality: position, opacity, scale, and rotation. Based on these bases, facial appearances are generated by a linear combination.
@@ 2.1 NeRF-based avatars
Numerous methods [91, 41, 64, 82, 86, 58, 11, 79, 25, 52, 34, 16, 92] capitalize on the speed and quality of 3DGS. Qian et al. [58] attach 3D Gaussians to the FLAME [40] mesh surface and apply a deformation gradient similar to Zielonka et al. [89] to orient the Gaussians according to the local Frenet frames of the surface. This method, however, does not utilize any information about expressions and, thus, struggles with pose-dependent changes (e.g., wrinkles, self-shadows) and, despite high-quality results, retrieves only a global static appearance model. 3D Gaussian blendshapes [46] controls an avatar by linearly interpolating between optimized blendshapes using 3DMM expression coefficents. However, this method depends on an underlying 3DMM whereas GEM is a mesh-free representation. Li et al. [41] use a StyleUNet-like CNN architecture [75] to regress front and back Gaussian maps. Employing a powerful CNN network on position maps, they achieve impressive results for human bodies with effects like pose-dependent wrinkle formation.
@@ 3.1 Gaussian Eigen Model (GEM)
We are following the idea of organizing the 3D Gaussians in 2D maps [41, 52, 82, 64], where each pixel represents a 3D Gaussian with its parameters. We propose an adapted CNN-architecture of Animatable Gaussians (AG) [41], by merging the separate Style-U-Nets, reducing the convolutional layers, and operating in the UV space of the FLAME head model. In addition, we are employing deformation gradients following Sumner et al. [65] to handle the transformation from canonical to deformed space and treat the color as a global parameter. We refer to the suppl. mat. for a detailed explanation of the architectural changes. In comparison to the original AG model, our proposed CNN model produces slightly better results while being more efficient in terms of computing and memory. Using this model, we generate the per-frame Gaussian primitives $\boldsymbol{G}_{i}$ in the canonical space for all training time-frames. Note that for this reconstruction, we follow Animatable Gaussians and, thus, FLAME-based tracking is required. However, during inference, our model is independent of FLAME.
@@ 4.2 Cross-person Reenactment Evaluation
Figure 8: Compression error depending on the number of used principal components in GEM. The heatmaps show the photometric $\ell_{1}$-error for 10, 30, and 50 components using $128^{2}$ Gaussian maps. See suppl. doc. for additional evaluations.
Drawing inspiration from EMOCA [7], we further assess cross-re-enactment quantitatively by leveraging emotion recognition feature vectors from both the source image and the resulting cross-re-enactment, utilizing EmoNet [74]. For each pair of input and output images, we predict EmoNet features and measure cosine distance and $\mathcal{L}_{1}$ error between them. We report the numbers in the Table 3. Additionally, we also report FID scores [21] and rendering speed. Our method achieves on-par quality with the CNN-based solution while maintaining the highest frame rates and outperforming GA in terms of quality.
@@ 4.3 GEM Ablation Studies
We are interested in the compression error introduced by the projection on different amounts of principal components used in GEM, also concerning the memory consumption.
@@
Our smallest model weighs as little as 7MB using only 10 components of the eigenbasis. This is almost 12 times less than our smallest CNN-based model and almost 70 times less than Animatable Gaussians [41]. In contrast to neural networks, we can easily trade quality over size which is very useful in the context of different commodity devices with reduced compute capabilities. Table 4 presents how compression affects the quality of reconstruction, where we evaluate a sequence with $\sim 1$k frames for a single actor under a novel view. As expected, using only 10 components impacts the quality the most, however, the results are still of high quality, see Figure 8. Gaussian Avatars [58] offers a small size of the stored Gaussians cloud, ranging from 5MB, and 14MB without the FLAME model for $128^{2}$, $256^{2}$ Gaussians, respectively. However, the quality of reconstruction lacks wrinkle details and sharpness as can be seen in Figure 10. In comparison to Gaussian Avatars [58], our model does not require FLAME during inference which is an additional 90MB.
@@
Figure 9 demonstrates that our method is able to handle different topologies (subject wearing glasses), despite utilizing a fixed UV space.
@@ B.2 CNN Training Details
Figure 13: The quality comparison to Gaussian Avatars [58] shows better performance, even though our model is similar in size to its Gaussian cloud, and we do not need an additional FLAME model, which weighs 90 MB.